# thanks: Corresponding author

### v2

@@ Abstract
In the framework of mutually unbiased bases (MUBs), a measurement in one basis gives no information about the outcomes of measurements in another basis. Here, we relax the no-information condition by allowing the $d$ outcomes to be predicted according to a predefined probability distribution $q=(q_{0},\ldots,q_{d-1})$. The notion of mutual unbiasedness, however, is preserved by requiring that the extracted information is the same for any preparation and any measurement; regardless of which state from which basis is chosen to prepare the system, the outcomes of measuring the system with respect to the other basis generate the same probability distribution. In the light of this, we define the notion of mutually equi-biased bases (MEBs) such that within each basis the states are equi-biased with respect to the states of the other basis and that the bases are mutually equi-biased with respect to each other. For $d=2,3$, we derive a complete set of $d+1$ MEBs. The mutual equi-biasedness imposes nontrivial constraints on the distribution $q$, leading for $d=3$ to the restriction $1/3\leq\mu\leq 1/2$ where $\mu=\sum_{k=0}^{2}q_{k}^{2}$. To capture the incompatibility of the measurements in MEBs, we derive an inequality for the probabilities of projective measurements in a qudit system, which yields an associated entropic uncertainty inequality. Finally, we construct a class of positive maps and their associated entanglement witnesses based on MEBs. While an entanglement witness constructed from MUBs is generally finer than one based on MEBs when both use the same number of bases, for certain values of the index $\mu$, employing a larger set of MEBs can yield a finer witness. We illustrate this behavior using isotropic states of a $3\times 3$ system. Our results reveal that not all bases in a set of $L$ MEBs can have contribution for the entanglement detection. A constraint, dependent on the probability distribution $q$, limits the number of MEBs that can be employed.
mutually unbiased bases; mutually equi-biased bases; entropic uncertainty relation; entanglement witness
@@ I Introduction
By generalizing the notion of unbiasedness from projective measurements to more general positive operator-valued measures (POVMs), Kalev et al. KalevNJP2014 introduced the notion of mutually unbiased measurements (MUMs), which encompass a broader class of measurements that retain the essential features of unbiasedness. Mutually unbiased bases and measurements have found important applications in quantum information, serving both as effective methods for entanglement detection SpenglerPRA2012 ; ChruscinskiPRA2014 ; ChenPRA2014 ; ChruscinskiPRA2018 ; LiIJTP2019 ; SalehiQIP2021 , and as foundational frameworks for entropic certainty and uncertainty relations MaassenPRL1988 ; ColesRMP2017 ; BallesterPRA2007 ; MolmerPRA2009 ; WehnerNJP2010 ; SanchezPLA1993 ; ZyczkowskiPRA2015 ; HuangPRA2024 ; HuangPRResearch2024 . Whereas entropic uncertainty relations set lower bounds on average entropies, reflecting the impossibility of a state being a simultaneous eigenstate of all observables, entropic certainty relations set upper bounds, arising from the fact that—apart from the completely mixed state—no quantum state can produce $d$ equally likely outcomes for every observable at once SanchezPLA1993 . For the average entropy characterizing measurements of a pure quantum state of dimension $d$ in $L$ orthogonal bases, both lower and upper bounds have been established in ZyczkowskiPRA2015 .
Here we extend this notion of MUBs to define mutually equi-biased bases (MEBs) such that the standard MUBs are obtained as a special limiting case of our results. To this aim, the unbiasedness condition is partially relaxed for MEBs in the sense that the transition probability from any state of one basis to any state of the second basis is no longer necessarily equal to $1/d$. In the most general setting of quantum measurement, if a system is prepared in a state belonging to one basis, the statistics of the outcomes of the measuring the system with respect to the states of the other basis provides a probability distribution ${q}=(q_{0},\ldots,q_{d-1})$. We adapt this notion of biasedness, however, our restriction is that regardless of which state of the first basis is chosen, the probability distribution of the transition to each state of the second basis remains the same up to a permutation of its entries. Accordingly, within each basis the states are equi-biased with respect to the states of the other basis and that the bases are mutually equi-biased with respect to each other. In the light of this, we associate a probability distribution ${q}=(q_{0},\ldots,q_{d-1})$ with each of MEBs such that
@@
In MEBs, in contrary to MUBs, the knowledge of the measuring the system with respect to the one basis, when the system is prepared in any state of the other basis, leaks some information about the state of the system. The mutual equi-biasedness of the bases implies, however, that they are biased in the same degree in the sense that regardless of which state from which basis are chosen to prepare the system, the extracted information is the same, i.e., they are predictable in the same degree with respect to each other. It makes sense therefore that the question of how much information can be extracted should be related to how far the probability distribution is from a homogenous distribution. The index of coincidence, defined by $\mu(q)=\sum_{k=0}^{d-1}q_{k}^{2}$ for the probability $q=(q_{0},\ldots,q_{d-1})$, is a measure of how much a probability distribution deviates from a uniform distribution, as such, it can play a key role when the extracted information is quantified. Hereafter, for brevity, we will refer to the index of coincidence simply as the index and denote it by $\mu$ for a $d$-outcomes probability distribution $q=(q_{0},\ldots,q_{d-1})$.
Related to the index $\mu$, we now pose the question that whether for a given $d$, a complete set of $d+1$ MEBs exists for all values of $\mu\in[1/d,1]$? We derive a complete set of MEBs for two cases $d=2,3$. Our results show that except for the particular case of $2$-dimensional system for which a complete set of $3$ MEBs exists for all $\mu\in[1/2,1]$, it is not the case for $d\geq 3$. More precisely, for $d=3$, the constraints imposed by the mutual equi-biasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists.
@@
The paper is organized as follows. In Sec. II, the Preliminaries subsection briefly reviews the relevant concepts of bistochastic and unistochastic matrices, along with a parameterization for probability distribution ${q}=(q_{0},\ldots,q_{d-1})$. We then introduce the notion of MEBs and present a complete derivation of MEBs for the two simplest cases, $d=2,3$. In Sec. III, we derive an inequality for the probabilities of projective measurements in MEBs of a qudit system. Based on this, we provide an entropic uncertainty inequality in MEBs. In Sec. IV, we construct a positive map based on MEBs, then introduce an entanglement witness and examine its performance for two-qutrit isotropic states. Finally, we conclude the paper in Sec. V with a brief summary. Some proofs of the results are provided in Appendix A.
II Mutually equi-biased bases (MEBs)
@@ II.1.1 Bistochastic and unistochastic matrices
We begin this section by defining the notion of mutually equi-biased bases (MEBs), which generalizes the concept of mutually unbiased bases (MUBs) to encompass a broader family of probability distributions. We will see that the allowed region of this family is shaped by the mutually equi-biased condition.
@@
Let $q=(q_{0},\ldots,q_{d-1})$ be a probability distribution over $d$ outcomes. We say that the set of $L$ orthonormal bases $\{\mathcal{B}^{(1)},\ldots,\mathcal{B}^{(L)}\}$ of the $d$-dimensional Hilbert space $\mathcal{H}$ are MEBs associated to the probability distribution $q$, if for every pair $\alpha\neq\beta$, the corresponding projection operators satisfy Eq. (2) for some permutation $\sigma$. Here, $\mathcal{B}^{(\alpha)}=\{|e_{k}^{(\alpha)}\rangle\}$ where $\alpha=1,\ldots,L$, and $k=0,\ldots,d-1$.
@@
More precisely, the set of matrices $\mathcal{Q}$s forms a (d1)(d-1)-dimensional subclass of the bistochastic matrices of size $d\times d$, which themselves comprise a convex set of dimension $(d-1)^{2}$, known as Birkhoff’s polytope ZyczkowskiCMP2005 . A bistochastic matrix $\mathcal{Q}$ is called unistochastic matrix if and only if its entries are the squares of the absolute values of the entries of some unitary matrix. In addition to the normalization requirement—i.e., in each row and column of a unitary matrix, the squared moduli sum to unity—the moduli of its entries must also satisfy certain constraints arising from the mutual orthogonality of the rows and columns. Accordingly, not all bistochastic matrices arise from unitary matrices; nonetheless, the $d\times d$ bistochastic matrix $\mathcal{Q}$, associated with the transition probabilities between two orthonormal bases $\mathcal{B}^{(\alpha)}$ and $\mathcal{B}^{(\beta)}$ is a unistochastic matrix. The characterization of the unistochastic subset of Birkhoff’s polytope, i.e., those for which there exist a unitary matrix $\mathcal{U}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, is considered in ZyczkowskiCMP2005 .
@@
To set the convention, suppose $\mathcal{B}^{(1)}=\{|e_{k}^{(1)}\rangle\}_{k=0}^{d-1}$ denotes the standard (computational) basis of the $d$-dimensional Hilbert space $\mathcal{H}$. Then, for each $\beta\in\{2,\ldots,L\}$, the orthonormal basis $\mathcal{B}^{(\beta)}$ is related to $\mathcal{B}^{(1)}$ via a unitary matrix $\mathcal{U}^{(\beta)}$ as $|e_{k}^{(\beta)}\rangle=\mathcal{U}^{(\beta)}|e_{k}^{(1)}\rangle$, for $k=0,\ldots,d-1$, where $\mathcal{U}^{(\beta)}_{kl}=\langle e_{k}^{(1)}|e_{l}^{(\beta)}\rangle$. By equi-biasedness, we have to assign $|\mathcal{U}_{kl}^{(\beta)}|^{2}=\mathcal{Q}_{kl}$, for all $\beta\in\{2,\ldots,L\}$, where the $d\times d$ bistochastic matrix $\mathcal{Q}$ is defined by $\mathcal{Q}_{kl}=q_{k\oplus l}$ for the probability distribution $q=(q_{0},\ldots,q_{d-1})$. An important question is to determine for which probability distributions $q$ is the associated bistochastic matrix $\mathcal{Q}$ unistochastic—that is, which $q$ gives rise to unistochastic matrix $\mathcal{Q}$ corresponding to unitary matrices $\mathcal{U}^{(\beta)}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$?
@@
Before we proceed further, let us first note that for a given unistochastic matrix $\mathcal{Q}$, if one unitary matrix $\mathcal{U}$ is found such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, then there exist infinitely many unitary matrices that satisfy this condition. This follows from the fact that the left and right multiplication of a unitary matrix $\mathcal{U}$ by any diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ results in another unitary matrix $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$, but the associated unistochastic matrix remains unchanged. This defines an equivalence relation as $\mathcal{U}\sim\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ ZyczkowskiJMP2009 . Note, however, that the equivalence is partial in general; not any pair of unitary matrices $\mathcal{U}$ and $\mathcal{V}$ generating the same unistochastic matrix $\mathcal{Q}$ can be written as $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ for some diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. For $d=2$, the correspondence between a unitary matrix and its unistochastic image is one-to-one up to the multiplications from the left and right by diagonal unitary matrices. For higher dimensions, the equivalence is not complete, meaning that, inequivalent unitaries can generate the same unistochastic matrix. An example occurs in dimension $d=4$ for a one-parameter class of complex Hadamard matrices ZyczkowskiOSID2006 . For $d=3$, to the best of our knowledge, there is no general proof that inequivalent unitaries cannot exist, nor any counterexample indicating that such inequivalent unitaries do exist.
@@
What is crucial in our analysis, however, is the requirement of the mutual equi-biasedness. This condition imposes a severe restriction on the set of MEBs, making it finite and constrained. To be more precise, suppose $\mathcal{Q}$ is a unistochastic matrix generated by L1L-1 unitary matrices $\mathcal{U}^{(\beta)}$, such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$, for each $\beta\in\{2,\ldots,L\}$. The mutually equi-biasedness requires further that they are mutually unistochastic, i.e., $|[{\mathcal{U}^{(\alpha)}}^{\dagger}\mathcal{U}^{(\beta)}]_{kl}|^{2}=\mathcal{Q}_{\sigma(k)\sigma(l)}$ for all $\alpha,\beta\in\{2,\ldots,L\}$ with $\alpha\neq\beta$. The permutation $\sigma$, appearing here and in Eq. (2), depends on $\alpha$ and $\beta$ in general, however, when one of the basis corresponds to the standard basis, we follow the convention that the permutation $\sigma$ does nothing, i.e., $\sigma(k)=k$ and $\sigma(l)=l$.

### v3

@@ Abstract
In the framework of mutually unbiased bases (MUBs), a measurement in one basis gives no information about the outcomes of measurements in another basis. Here, we relax the no-information condition by allowing the $d$ outcomes to be predicted according to a predefined probability distribution $q=(q_{0},\ldots,q_{d-1})$. The notion of mutual unbiasedness, however, is preserved by requiring that the extracted information is the same for any preparation and any measurement; regardless of which state from which basis is chosen to prepare the system, the outcomes of measuring the system with respect to the other basis generate the same probability distribution. In light of this, we define the notion of mutually equibiased bases (MEBs) such that within each basis the states are equibiased with respect to the states of the other basis and that the bases are mutually equibiased with respect to each other. For $d=2,3$, we derive a set of $d+1$ MEBs. The mutual equibiasedness imposes nontrivial constraints on the distribution $q$, leading for $d=3$ to the restriction $1/3\leq\mu\leq 1/2$ where $\mu=\sum_{k=0}^{2}q_{k}^{2}$. To capture the incompatibility of the measurements in MEBs, we derive an inequality for the probabilities of projective measurements in a qudit system, which yields an associated entropic uncertainty inequality. Finally, we construct a class of positive maps and their associated entanglement witnesses based on MEBs. While an entanglement witness constructed from MUBs is generally finer than one based on MEBs when both use the same number of bases, for certain values of the index $\mu$, employing a larger set of MEBs can yield a finer witness. We illustrate this behavior using isotropic states of a $3\times 3$ system. Our results reveal that not all bases in a set of $L$ MEBs can contribute to the entanglement detection. A constraint, dependent on the probability distribution $q$, limits the number of MEBs that can be employed.
mutually unbiased bases; mutually equibiased bases; entropic uncertainty relation; entanglement witness
@@ I Introduction
By generalizing the notion of unbiasedness from projective measurements to more general positive operator-valued measures, Kalev and Gour KalevNJP2014 introduced the notion of mutually unbiased measurements, which encompass a broader class of measurements that retain the essential features of unbiasedness. Mutually unbiased bases and measurements have found important applications in quantum information, serving both as effective methods for entanglement detection SpenglerPRA2012 ; ChruscinskiPRA2014 ; ChenPRA2014 ; ChruscinskiPRA2018 ; LiIJTP2019 ; SalehiQIP2021 , and as foundational frameworks for entropic certainty and uncertainty relations MaassenPRL1988 ; ColesRMP2017 ; BallesterPRA2007 ; MolmerPRA2009 ; WehnerNJP2010 ; SanchezPLA1993 ; ZyczkowskiPRA2015 ; HuangPRA2024 ; HuangPRResearch2024 . Whereas entropic uncertainty relations set lower bounds on average entropies, reflecting the impossibility of a state being a simultaneous eigenstate of all observables, entropic certainty relations set upper bounds, arising from the fact that, apart from the completely mixed state, no quantum state can produce $d$ equally likely outcomes for every observable at once SanchezPLA1993 . For the average entropy characterizing measurements of a pure quantum state of dimension $d$ in $L$ orthogonal bases, both lower and upper bounds have been established in ZyczkowskiPRA2015 .
Here we extend this notion of MUBs to define mutually equibiased bases (MEBs) such that the standard MUBs are obtained as a special limiting case of our results. To this aim, the unbiasedness condition is partially relaxed for MEBs in the sense that the transition probability from any state of one basis to any state of the second basis is no longer necessarily equal to $1/d$. In the most general setting of quantum measurement, if a system is prepared in a state belonging to one basis, the statistics of the outcomes of measuring the system with respect to the states of the other basis provides a probability distribution ${q}=(q_{0},\ldots,q_{d-1})$. We adapt this notion of biasedness; however, our restriction is that regardless of which state of the first basis is chosen, the probability distribution of the transition to each state of the second basis remains the same up to a permutation of its entries. Accordingly, within each basis the states are equibiased with respect to the states of the other basis and that the bases are mutually equibiased with respect to each other. In light of this, we associate a probability distribution ${q}=(q_{0},\ldots,q_{d-1})$ with each of MEBs such that
@@
In MEBs, in contrast to MUBs, the knowledge of measuring the system with respect to the one basis, when the system is prepared in any state of the other basis, leaks some information about the state of the system. The mutual equibiasedness of the bases implies, however, that they are biased in the same degree in the sense that regardless of which state from which basis is chosen to prepare the system, the extracted information is the same, i.e., they are predictable in the same degree with respect to each other. It therefore makes sense that the question of how much information can be extracted should be related to how far the probability distribution is from a homogenous distribution. The index of coincidence, defined by $\mu(q)=\sum_{k=0}^{d-1}q_{k}^{2}$ for the probability $q=(q_{0},\ldots,q_{d-1})$, is a measure of how much a probability distribution deviates from a uniform distribution; as such, it can play a key role when the extracted information is quantified. Hereafter, for brevity, we will refer to the index of coincidence simply as the index and denote it by $\mu$ for a $d$-outcome probability distribution $q=(q_{0},\ldots,q_{d-1})$.
Related to the index $\mu$, we now pose the question of whether, for a given $d$, a set of $d+1$ MEBs exists for all values of $\mu\in[1/d,1]$. We derive a set of $d+1$ MEBs for two cases $d=2,3$. Our results show that except for the particular case of two-dimensional system for which a set of three MEBs exists for all $\mu\in[1/2,1]$, it is not the case for $d\geq 3$. More precisely, for $d=3$, the constraints imposed by the mutual equibiasedness limits the range of the index $\mu$ to the feasible region $1/3\leq\mu\leq 1/2$ within which a set of four MEBs exists.
@@
The paper is organized as follows. In Sec. II we briefly review the relevant concepts of bistochastic and unistochastic matrices, along with a parameterization for probability distribution ${q}=(q_{0},\ldots,q_{d-1})$. We then introduce the notion of MEBs and present a complete derivation of MEBs for the two simplest cases, $d=2,3$. In Sec. III, we derive an inequality for the probabilities of projective measurements in MEBs of a qudit system. Based on this, we provide an entropic uncertainty inequality in MEBs. In Sec. IV, we construct a positive map based on MEBs and then introduce an entanglement witness and examine its performance for two-qutrit isotropic states. We conclude the paper in Sec. V with a brief summary. Some proofs of the results are provided in Appendices A and B.
II Mutually equibiased bases (MEBs)
@@ II.1.1 Bistochastic and unistochastic matrices
We begin this section by defining the notion of mutually equibiased bases, which generalizes the concept of mutually unbiased bases to encompass a broader family of probability distributions. We will see that the allowed region of this family is shaped by the mutually equibiased condition.
@@
Let $q=(q_{0},\ldots,q_{d-1})$ be a probability distribution over $d$ outcomes. We say that the set of $L$ orthonormal bases $\{\mathcal{B}^{(1)},\ldots,\mathcal{B}^{(L)}\}$ of the $d$-dimensional Hilbert space $\mathcal{H}$ is MEBs associated with the probability distribution $q$ if for every pair $\alpha\neq\beta$ the corresponding projection operators satisfy Eq. (2) for some permutation $\sigma$. Here $\mathcal{B}^{(\alpha)}=\{|e_{k}^{(\alpha)}\rangle\}$ where $\alpha=1,\ldots,L$ and $k=0,\ldots,d-1$.
@@
More precisely, the set of matrices $\mathcal{Q}$ forms a (d1)(d-1)-dimensional subclass of the bistochastic matrices of size $d\times d$, which themselves comprise a convex set of dimension $(d-1)^{2}$, known as Birkhoff’s polytope ZyczkowskiCMP2005 . A bistochastic matrix $\mathcal{Q}$ is called unistochastic matrix if and only if its entries are the squares of the absolute values of the entries of some unitary matrix. In addition to the normalization requirement, i.e., in each row and column of a unitary matrix the squared moduli sum to unity, the moduli of its entries must also satisfy certain constraints arising from the mutual orthogonality of the rows and columns. Accordingly, not all bistochastic matrices arise from unitary matrices; nonetheless, the $d\times d$ bistochastic matrix $\mathcal{Q}$, associated with the transition probabilities between two orthonormal bases $\mathcal{B}^{(\alpha)}$ and $\mathcal{B}^{(\beta)}$ is a unistochastic matrix. The characterization of the unistochastic subset of Birkhoff’s polytope, i.e., that for which there exist a unitary matrix $\mathcal{U}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, is considered in ZyczkowskiCMP2005 .
@@
To set the convention, suppose $\mathcal{B}^{(1)}=\{|e_{k}^{(1)}\rangle\}_{k=0}^{d-1}$ denotes the standard (computational) basis of the $d$-dimensional Hilbert space $\mathcal{H}$. Then, for each $\beta\in\{2,\ldots,L\}$, the orthonormal basis $\mathcal{B}^{(\beta)}$ is related to $\mathcal{B}^{(1)}$ via a unitary matrix $\mathcal{U}^{(\beta)}$ as $|e_{k}^{(\beta)}\rangle=\mathcal{U}^{(\beta)}|e_{k}^{(1)}\rangle$, for $k=0,\ldots,d-1$, where $\mathcal{U}^{(\beta)}_{kl}=\langle e_{k}^{(1)}|e_{l}^{(\beta)}\rangle$. By equibiasedness, we have to assign $|\mathcal{U}_{kl}^{(\beta)}|^{2}=\mathcal{Q}_{kl}$, for all $\beta\in\{2,\ldots,L\}$, where the $d\times d$ bistochastic matrix $\mathcal{Q}$ is defined by $\mathcal{Q}_{kl}=q_{k\oplus l}$ for the probability distribution $q=(q_{0},\ldots,q_{d-1})$. An important question is to determine for which probability distributions $q$ the associated bistochastic matrix $\mathcal{Q}$ is unistochastic, that is, which $q$ gives rise to unistochastic matrix $\mathcal{Q}$ corresponding to unitary matrices $\mathcal{U}^{(\beta)}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$.
@@
Before we proceed further, let us first note that for a given unistochastic matrix $\mathcal{Q}$, if one unitary matrix $\mathcal{U}$ is found such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}|^{2}$, then there exist infinitely many unitary matrices that satisfy this condition. This follows from the fact that the left and right multiplication of a unitary matrix $\mathcal{U}$ by any diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ results in another unitary matrix $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$, but the associated unistochastic matrix remains unchanged. This defines an equivalence relation as $\mathcal{U}\sim\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ ZyczkowskiJMP2009 . Note, however, that the equivalence is partial in general; not any pair of unitary matrices $\mathcal{U}$ and $\mathcal{V}$ generating the same unistochastic matrix $\mathcal{Q}$ can be written as $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ for some diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. For $d=2$, the correspondence between a unitary matrix and its unistochastic image is one to one up to the multiplications from the left and right by diagonal unitary matrices. For higher dimensions, the equivalence is not complete, meaning that inequivalent unitaries can generate the same unistochastic matrix. An example occurs in dimension $d=4$ for a one-parameter class of complex Hadamard matrices ZyczkowskiOSID2006 . For $d=3$, to the best of our knowledge, there is no general proof that inequivalent unitaries cannot exist, nor any counterexample indicating that such inequivalent unitaries do exist.
@@
What is crucial in our analysis, however, is the requirement of the mutual equibiasedness. This condition imposes a severe restriction on the set of MEBs, making it finite and constrained. To be more precise, suppose $\mathcal{Q}$ is a unistochastic matrix generated by L1L-1 unitary matrices $\mathcal{U}^{(\beta)}$ such that $\mathcal{Q}_{kl}=|\mathcal{U}_{kl}^{(\beta)}|^{2}$ for each $\beta\in\{2,\ldots,L\}$. The mutually equibiasedness requires further that they are mutually unistochastic, i.e., $|[{\mathcal{U}^{(\alpha)}}^{\dagger}\mathcal{U}^{(\beta)}]_{kl}|^{2}=\mathcal{Q}_{\sigma(k)\sigma(l)}$ for all $\alpha,\beta\in\{2,\ldots,L\}$ with $\alpha\neq\beta$. The permutation $\sigma$, appearing here and in Eq. (2), depends on $\alpha$ and $\beta$ in general, however, when one of the bases corresponds to the standard basis, we follow the convention that the permutation $\sigma$ does nothing, i.e., $\sigma(k)=k$ and $\sigma(l)=l$.